OpenStudy (anonymous):

Under the action of a constant force an object accelerates at 7.6 m/s^2. What will the acceleration be if (a) The force is halved ? (b) The object's mass is halved? (c) The force and the the object's mass are both halved? (d) The force is halved and the object's mass is doubled?

OpenStudy (cwrw238):

newtons second law force = mass * acceleration

OpenStudy (cwrw238):

i'll do (a) for you F = m * 7.6 0.5F = m * a the force is halved so right hand side of the latter equation must be halved so what does that make a?